OpenStudy (jacobbirddog):

Please Help! Algebra 1 quadratic formulas! Medal+fan! 2 Questions! What is the discriminant for the quadratic expression? 18x2 + 11x Use the discriminant to determine the number and type of solutions for the following equation. 24x2 + 13 = 0

ganeshie8 (ganeshie8):

\(\large ax^2 + bx+c\) discriminant = \(\large b^2 - 4ac\)

OpenStudy (jacobbirddog):

im thinking the actual numbers and they are looking like 121, but i think i am wrong

ganeshie8 (ganeshie8):

your quadratic : \(\large 18x^2 + 11x + 0\) \(\large a = 18\) \(\large b = 11\) \(\large c = 0\) discriminant = \(\large b^2-4ac \) \(\qquad = 11^2 - 4(18)(0)\) \(\qquad = 121\)

ganeshie8 (ganeshie8):

you're right !!

OpenStudy (jacobbirddog):

wow, im usually wrong in this type of maths,

ganeshie8 (ganeshie8):

lol, that wil change easily wid practice im sure :)

OpenStudy (jacobbirddog):

on the second one though, i do keep getting that one wrong in particular

OpenStudy (jacobbirddog):

not trying to leech :P

ganeshie8 (ganeshie8):

wat did u get for discriminant on second one ?

OpenStudy (jacobbirddog):

i got one rational solution

ganeshie8 (ganeshie8):

given quadratic : \(\large 24x^2 + 13 = 0\) \(\large a = 24\) \(\large b=0\) \(\large c = 13\) discriminant = ?

ganeshie8 (ganeshie8):

evaluate the discriminant first,

OpenStudy (jacobbirddog):

i got 15

OpenStudy (jacobbirddog):

i am sooo wrong...

OpenStudy (jacobbirddog):

let me redo the problem real quick

ganeshie8 (ganeshie8):

given quadratic : \(\large 24x^2 + 13 = 0\) \(\large a = 24\) \(\large b=0\) \(\large c = 13\) discriminant = \(\large b^2-4ac = 0^2 - 4(24)(13) = -1248\)

ganeshie8 (ganeshie8):

So, its negative discriminant !

ganeshie8 (ganeshie8):

what does a NEGATIVE discriminant tell u about the solutions ?